OpenStudy (lizilizi28):

muliply (x^2 + y^5)(x^2 + -y^5)

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):



OpenStudy (anonymous):

x^4-y^10

OpenStudy (anonymous):

if (a+b)*(c+d) = (x^2 + y^5)(x^2 + -y^5) a b c d then a*c+a*d+b*c+b*d

OpenStudy (anonymous):

make it simple by letting x^2 = a y^5 = b so you have: (a+b)(a-b) then use distribution method. after getting the answer, substitute the values to a and b.

jimthompson5910 (jim_thompson5910):

this is another difference of squares problem (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

didn't catch that lol but my way will still solve it